OpenStudy (anonymous):

Twin brothers, Billy and Bobby, can mow their grandparent's lawn in 76 minutes. Billy could mow the lawn by himself in 15 minutes less time than it would take Bobby. How long would it take Bobby to mow the lawn by himself?

OpenStudy (amistre64):

T(bob) + T(bill) = 76 is what im thinking

OpenStudy (amistre64):

bob-15 + bob = 76

OpenStudy (amistre64):

2bob = 91 bob = 91/2 = 45.5 maybe

OpenStudy (amistre64):

that almost feels right; almost ...

OpenStudy (amistre64):

the by himself part is not being taken into account

OpenStudy (anonymous):

amistre 64 ineed ur help

OpenStudy (amistre64):

I dont know c++

OpenStudy (anonymous):

oh....ok...:(

OpenStudy (amistre64):

if we had the set up the we know bob and bill alone we could get to time together .... so im gonna have to look backwards at this :)

OpenStudy (amistre64):

bob = time (b) alone bill = time (b -15) alone \[\frac{1}{b}+\frac{1}{b-15}=\frac{1}{76}\] \[\frac{b-15}{b(b-15)}+\frac{b}{b(b-15)}=\frac{1}{76}\] \[\frac{b-15+b}{b(b-15)}=\frac{1}{76}\] \[\frac{b(b-15)}{b-15+b}=76\] \[\frac{b^2-15b}{2b-15}=76\] \[b^2-15b=76(2b-15)\] \[b^2-15b-76(2b-15)=0\] \[b^2-15b-152b+1140=0\] \[b^2-167b+1140=0\] maybe?

OpenStudy (amistre64):

b = about 160 or 8; so id go with 160 minutes as a gut thing if:\[\frac{1}{160}+\frac{1}{160-15}=\frac{1}{76}\]then it should be good

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=1%2F160+%2B+1%2F%28160-15%29+-+1%2F76 yeah, im going with that, or close to it